OpenStudy (anonymous):

The equation of line CD is y = 3x − 3. Write an equation of a line perpendicular to line CD in slope-intercept form that contains point (3, 1). y = 3x + 0 y = −3x − 8 y = negative 1 over 3x + 2 y = − 1 over 3x + 0 @tester97

OpenStudy (jdoe0001):

hmmm what's the slope of => y = 3x − 3. ?

OpenStudy (anonymous):

3x?? @jdoe0001

OpenStudy (jdoe0001):

\(\bf \begin{array}{llll} y = &{\color{brown}{ 3}}x - 3\\ &\uparrow \\ &{\color{brown}{ slope}} \end{array}\) so the slope of a line perpendicular to this one, will have a slope that is, NEGATIVE RECIPROCAL to this slope so this one is 3 so \(\bf 3\qquad negative\to -3 \qquad reciprocal\to \cfrac{1}{-3}\) so what you're really asked is, to find the equation of a line whose slope is -1/3 and passes through (3,1) thus \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ 1}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -\frac{1}{3} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

solving for "y" will more or less give you the -> y = mx+b <-- form, so-called the slope-intercept form

OpenStudy (anonymous):

i still dont get it :( @jdoe0001

OpenStudy (jdoe0001):

ok... what part?

OpenStudy (anonymous):

the plug in numbers.. @jdoe0001

OpenStudy (jdoe0001):

well, notice the values, they're pretty much given

OpenStudy (anonymous):

sooo it would be c?? @jdoe0001